OpenStudy (anonymous):

need to round improper fraction 149/12 to the nearest tenth

OpenStudy (anonymous):

This is\[ \frac{149}{12}=12\frac{5}{12} \]We know: \[ 0.4<\frac{5}{12}<0.45\implies\frac{5}{12}\approx0.4 \]So: \[ 12+0.4=12.4 \]You could also divide, which is easier, but I like multiplying, better. Enjoy.